# Sharpness of Minima in Deep Matrix Factorization: Exact Expressions

### v2

@@ C.1 Visualization of the Contour Map of the Loss Landscape
C.2 Additional Experiments
For the experiment, we first generate the layer dimensions randomly and then construct the optimal layers $\bm{W}_{1}^{,}\bm{W}_{2}^{,}\dots,\bm{W}_{L}^{*}]$. We then perform the same experiments as in Figs. 13, varying the depth, dimensions, and initialization distance $r$ (as shown in Figs. 17). We note that oscillations occur along the eigenvector corresponding to the maximum eigenvalue of the Hessian. The dimensions of the factors, i.e., $d_{0},d_{1},\dots,d_{L}$}, in Fig. 6 are given by 1,9,4,8,24,16,17,11,21,3,22,3,3,15,3,18,17,16,5,12,1{1,9,4,8,24,16,17,11,21,3,22,3,3,15,3,18,17,16,5,12,1}, which implies $N=2421$, while the dimensions of the factors in Fig. 7 are given by 1,9,4,8,1{1,9,4,8,1}, which implies $N=293$.

### v3

@@ D.1 Visualization of the Contour Map of the Loss Landscape
D.2 Experiment Details and Additional Experiments
Figure 3: Contour map of the loss landscape around a minimum in a 15-layer overparameterized scalar factorization of a random scalar. GD with different step sizes $\eta$, indicated by different colors, is initialized within a radius of $10^{-9}$ from the minimum, in the direction of the Hessian eigenvector corresponding to the largest eigenvalue. The vector $\bm{v}_{1}$ denotes the eigenvector of the Hessian corresponding to the largest eigenvalue, while $\bm{v}_{N}$ denotes the eigenvector corresponding to the smallest eigenvalue. The value of $\lambda_{\max}(\nabla^{2}\mathcal{L}(\bm{w}^{*}))$ is computed using the closed-form expression derived in Theorem 21.
For the experiment, we first generate the layer dimensions randomly and then construct the optimal layers $\bm{W}_{1}^{*},\bm{W}_{2}^{*},\dots,\bm{W}_{L}^{*}]$ as Gaussian random matrices, with each entry sampled from $N(0,1)$ according to the generated dimensions. Then, we compute $\bm{M}$ or $m$ by $\prod_{j=1}^{L}\bm{W}^{*}_{j}$. We then perform the same experiments as in Figs. 12, varying the depth, dimensions, and initialization distance $r$ (as shown in Figs. 3-7). We note that oscillations occur along the eigenvector corresponding to the maximum eigenvalue of the Hessian. The dimensions of the factors, i.e., $d_{0},d_{1},\dots,d_{L}$}, in Fig. 6 are given by 1,9,4,8,24,16,17,11,21,3,22,3,3,15,3,18,17,16,5,12,1{1,9,4,8,24,16,17,11,21,3,22,3,3,15,3,18,17,16,5,12,1}, which implies $N=2421$, while the dimensions of the factors in Fig. 7 are given by 1,9,4,8,1{1,9,4,8,1}, which implies $N=293$.